# On the Exteneded Painlevé Property of Riccati Hiearchies

### v1

@@ 5 Extended Painlevé Property
$\left(X_{1},\dots,X_{m},Z,\varepsilon\right)=\begin{cases}\left(\frac{(n)_{1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-1/j},\dots,\frac{(n)_{j-1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-(j-1)},0,\dots,0,0\right),&\text{for }j\leq n,\\ (0,\dots,0),&\text{for }j>n,\end{cases}$ (5.3)
is a fixed point of the vector field in (5.2).
Lemma 5.2.
The eigenvalues of the Jacobi matrix of the vector field (5.2) at the fixed point (5.3) are given by $r$, $s$, and the Kovalevskaya exponents of the system (4.1), excluding the trivial exponent 1-1. This result has been established in [5].
We now state the main result of this section:
@@
The Riccati chain satisfies the Extended Painlevé Property (EPP) in the sense that it admits an mk+1m-k+1-parameter family of convergent Laurent series solutions, where the set $\{a_{i,n}\}$ includes mkm-k free parameters.
@@
For each indicial locus $(\xi^{(n)})$, it corresponds to n1n-1 negative eigenvalues, mnm-n positive eigenvalues, and the trivial exponent 1-1. By Lemma 4.2, the eigenvalues of the Jacobi matrix at the fixed point are given by $\kappa$, and the parameters $r$ and $s$.
By translation, the system can be decomposed into an mn+2m-n+2-dimensional unstable manifold and an n1n-1-dimensional stable manifold. We perform a suitable linear transformation, which allows us to rewrite the system in the form:
$\left\{\begin{array}[]{l}\displaystyle\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{u}\in\mathbb{C}^{m-n+2},\\ \displaystyle\frac{d\mbox{$X$}_{s}}{dt}=J_{s}\mbox{$X$}_{s}+F_{s}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{s}\in\mathbb{C}^{n-1}.\end{array}\right$.
where $J_{u}$ and $J_{s}$ are the Jacobian matrices corresponding to the unstable and stable manifolds, respectively, with the real parts of the eigenvalues of $J_{u}$ being positive and those of $J_{s}$ being negative.
By the unstable manifold theorem, there exists a local analytic function $\varphi$ such that $\varphi(0)=D\varphi(0)=0$, and the set $(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u}))$ forms the unstable manifold. The dynamics on the unstable manifold can then be expressed as:
$\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u})).$
Note that $J_{u}$ is semisimple and has distinct positive integer eigenvalues. Since for $j>n$, the fixed point is located at the origin, $(X_{1},\dots,X_{m},Z,\varepsilon)=(0,\dots,0,0)$, the fixed point lies at the origin.
Furthermore, since the system does not contain resonant terms in the Poincaré domain, we can apply Poincaré’s linearization theory to linearize the unstable manifold, completing the proof that the Riccati chain satisfies the Extended Painlevé Property. ∎

### v3

@@ 4 Commuting Vector Fields
$F,G]=0,\quad\text{where}\quad[F,G]_{i}=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{i}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{i}}{\partial x_{k}}(x)\right),\quad i=1,\dots,m.$
(A3) $F(x)=0$ only when $x=0$.
In the context of this paper, we consider the following proposition proposed in [16] related to the commutator and the indicial locus:
Proposition 4.1.
For a given indicial locus $c$, the identity $(K(c)+\gamma)G(c)=0$ holds. In particular, if $\gamma$ is not a K-exponent, then $G(c)=0$.
We now consider the autonomous Riccati chain $F$ on $\mathbb{C}^{m}$, where each component $f_{k}$ is given by:
$f_{k}(x)=x_{k+1}-cx_{1}x_{k},\quad 1\leq k\leq m-1,\quad f_{m}(x)=-cx_{1}x_{m},$
with $c\in\mathbb{C}^{*}$ and $x=(x_{1},\dots,x_{m})\in\mathbb{C}^{m}$. The weights of the variables are assumed to be $\mathrm{wt}(x_{k})=k$ for $k=1,\dots,m$.
Unique commuting vector field
@@
Let $F$ be the autonomous Riccati chain with weights $\mathrm{wt}(x_{k})=k$. The only nontrivial quasi-homogeneous polynomial vector field commuting with $F$ is the unique commuting field $G$ of degree $\gamma=m$, given by
$g_{i}=x_{i}x_{m}\quad\text{for}\quad i<m,\quad g_{m}=x_{m}^{2}.$
For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists.
@@
We start by considering the quasi-homogeneous vector field $G=(g_{1},g_{2},\dots,g_{m})$, where each component $g_{i}$ is a polynomial of degree $i+\gamma$, which can be written in the general form:
$g_{i}(x)=\sum_{j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma}c_{j_{1},j_{2},\dots,j_{m}}x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{m}^{j_{m}},$
where $c_{j_{1},j_{2},\dots,j_{m}}$ are constants, and the sum is over all multi-indices $(j_{1},j_{2},\dots,j_{m})$ such that the degree constraint $j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma$ is satisfied.
The commutator $F,G]_{m}(x)$ is given by:
$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{m}}{\partial x_{k}}(x)\right).$
Substituting the expressions for $f_{k}(x)$ and $g_{k}(x)$, we have:
$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1})\right).$
This simplifies to:
$F,G]_{m}(x)=\sum_{k=1}^{m-1}(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}(x)-cx_{1}x_{m}\frac{\partial g_{m}}{\partial x_{m}}-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1}).$
Now, we examine the critical term $(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}$. This term contains the factor $(x_{k}-cx_{1}x_{k-1})$, which introduces additional terms that cannot be canceled by any other terms in the expression. Specifically, when $k=m$, the structure of $g_{m}(x)=x_{m}^{2}$ creates terms that lead to an imbalance.
Thus, we find that the commutator $F,G]_{m}(x)$ cannot vanish unless $g_{i}(x)=x_{i}x_{m}$ for $i=1,\dots,m-1$ and $g_{m}(x)=x_{m}^{2}$, which is the only solution that satisfies the commutation relation.
Therefore, the only quasi-homogeneous polynomial vector fields commuting with $F$ are $F$ itself (degree 1) and the unique commuting field $G$ of degree $\gamma=m$, given by:
$g_{i}(x)=x_{i}x_{m}\quad\text{for}\quad i=1,\dots,m-1,\quad g_{m}(x)=x_{m}^{2}.$
For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists. ∎
Symmetry reduction
The commuting vector field $G$ simplifies the Riccati chain by generating invariants along its flow. Specifically, along the flow of $G$, the quantities $y_{i}=\frac{x_{i}}{x_{m}}$ remain constant, which means the ratios of the $x_{i}$’s relative to $x_{m}$ are invariant. These invariants, interpreted as symmetries of the system, allow for a reduction in the dimensionality of the system. By exploiting these invariants, we reduce the Riccati chain to a simpler triangular system, where the equations are decoupled and can be solved sequentially. The general solution is then obtained in terms of rational functions of an arbitrary polynomial $u(z)$, where the solutions for $x_{1},x_{2},\dots,x_{m}$ are expressed as functions of $u(z)$. For the rest of section, we assume $c=1$ without loss of generality.
Corollary 4.3.
Let $F$ be the autonomous Riccati chain (3.1) with $c=1$, i.e.
$x_{i}^{\prime}$ $=x_{i+1}-x_{1}x_{i},\qquad$ $1\leq i\leq m-1,$ (4.1)
$x_{m}^{\prime}$ $=-x_{1}x_{m}.$
Let $G$ be the commuting vector field of degree $\gamma=m$ given in Theorem 4.2,
$g_{i}=x_{i}x_{m}\ (i<m),\qquad g_{m}=x_{m}^{2}.$
On the open set $\{x_{m}\neq 0\}$, define
$y_{i}:=\frac{x_{i}}{x_{m}}\quad(i=1,\dots,m-1),\qquad u:=\frac{1}{x_{m}}.$ (4.2)
Then the Riccati chain (4.1) is transformed into the triangular system
$y_{i}^{\prime}=y_{i+1}\quad(1\leq i\leq m-2),\qquad y_{m-1}^{\prime}=1,\qquad u^{\prime}=y_{1}.$ (4.3)
Consequently, writing $t=z-z_{0}$ and introducing constants $a_{0},\dots,a_{m-1}\in{\mathbb{C}}$, one has
$u(t)=\frac{t^{m}}{m!}+a_{m-1}\frac{t^{m-1}}{(m-1)!}+\cdots+a_{1}t+a_{0},$ (4.4)
and
$y_{i}(t)=u^{(i)}(t)\qquad(i=1,\dots,m-1).$ (4.5)
The general solution of (4.1) is therefore
$x_{m}(z)=\frac{1}{u(z-z_{0})},\qquad x_{i}(z)=\frac{y_{i}(z-z_{0})}{u(z-z_{0})}=\frac{u^{(i)}(z-z_{0})}{u(z-z_{0})}\quad(i=1,\dots,m-1).$ (4.6)
Proof.
Define $y_{i},u$ by (4.2). For $1\leq i\leq m-1$,
$y_{i}^{\prime}=\left(\frac{x_{i}}{x_{m}}\right)^{\prime}=\frac{x_{i}^{\prime}x_{m}-x_{i}x_{m}^{\prime}}{x_{m}^{2}}.$
Using (4.1), for $1\leq i\leq m-2$ we obtain
$y_{i}^{\prime}=\frac{(x_{i+1}-x_{1}x_{i})x_{m}-x_{i}(-x_{1}x_{m})}{x_{m}^{2}}=\frac{x_{i+1}}{x_{m}}=y_{i+1},$
and for i=m1i=m-1,
$y_{m-1}^{\prime}=\frac{x_{m}}{x_{m}}=1.$
Next,
$u^{\prime}=\left(\frac{1}{x_{m}}\right)^{\prime}=-\frac{x_{m}^{\prime}}{x_{m}^{2}}=-\frac{-x_{1}x_{m}}{x_{m}^{2}}=\frac{x_{1}}{x_{m}}=y_{1}.$
This proves (4.3).
Integrating (4.3) yields $y_{m-1}(t)=t+a_{m-1}$, then $y_{m-2}(t)=\frac{t^{2}}{2!}+a_{m-1}t+a_{m-2}$, and inductively $y_{1}$ is a polynomial of degree m1m-1 whose coefficients are affine functions of the constants $a_{0},\dots,a_{m-1}$.
$y_{m-k}(t)=\frac{t^{k}}{k!}+\frac{a_{m-1}t^{k-1}}{(k-1)!}+\frac{a_{m-2}t^{k-2}}{(k-2)!}+\cdots+a_{m-k},\quad k=1,\dots,m-1.$
Finally, $u^{\prime}=y_{1}$ implies that $u$ is a polynomial of degree $m$, given by
$u(t)=\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}.$
The identities $y_{i}=u^{(i)}$ follow by differentiating $u^{\prime}=y_{1}$ repeatedly, hence (4.6) follows from $x_{m}=1/u$ and $x_{i}=y_{i}/u$. ∎
In particulae, the general solution $x_{1}(z)$ is written as
$x_{1}(z)$ $=\frac{y_{1}(z-z_{0})}{u(z-z_{0})}$ (4.7)
$=\frac{u^{(1)}(z-z_{0})}{u(z-z_{0})}$
$=\frac{\frac{t^{m-1}}{m-1!}+\frac{a_{m-1}t^{m-2}}{(m-2)!}+\cdots+a_{1}}{\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}}$
$=\frac{mt^{m-1}+m(m-1)a_{m-1}t^{m-2}+\cdots+m!a_{1}}{t^{m}+ma_{m-1}t^{m-1}+\cdots+m!a_{0}}$
$=\frac{mt^{m-1}+(m-1)b_{m-1}t^{m-2}+\cdots+(m-1)!b_{1}}{t^{m}+b_{m-1}t^{m-1}+\cdots+b_{0}}$
where $b_{m-k}=m(m-1)\cdots(m-k+1)a_{m-k}$.
Blow-up method
In the context of the 2-dimensional Riccati chain, we apply the blow-up method commonly used in Painlevé analysis to resolve singularities[13, 14]. This procedure provides a regular foliation on a smooth model and will be shown to yield the same general solution as derived through symmetry reduction. The blow-up method introduces a sequence of transformations that compactify the space and resolve indeterminacies at infinity, leading to a simpler structure that can be analyzed more easily.
Compactification and a base point at infinity
Consider the $2$-dimensional autonomous system
$x_{1}^{\prime}=x_{2}-x_{1}^{2},\qquad x_{2}^{\prime}=-x_{1}x_{2}.$ (4.8)
On $\mathbb{P}$$\times{\mathbb{P}}$ introduce the chart at infinity
$(X,Y):=\Bigl(\frac{1}{x_{1}},\frac{1}{x_{2}}\Bigr),$
so that the induced vector field is
$X^{\prime}=1-\frac{X^{2}}{Y},\qquad Y^{\prime}=\frac{Y}{X}.$
The point $(X,Y)=(0,0)$ is a base point, since the right-hand sides are indeterminate there.
Blow-up
Blow up $(X,Y)=(0,0)$. In the chart X=u,Y=uvX=u,\ Y=uv one obtains
$u^{\prime}=\frac{v-u}{v},\qquad v^{\prime}=1.$ (4.9)
Equation (4.9) is integrable:
$v(z)=z-z_{0},\qquad u(z)=\frac{(z-z_{0})^{2}+2C}{2(z-z_{0})}=\frac{z^{2}-2zz_{0}+z_{0}^{2}+2C}{2z-2z_{0}},$
where $C$ is an arbitrary constant.
With $b_{1}=2z_{0}$ and $b_{2}=z_{0}^{2}+2C$, we recover the same general solution as (4.7) after transforming back to $(x_{1},x_{2})$.
5 Annular Laurent series expansions
In this section, we examine the Laurent expansion of the solution $x_{1}(z)$ in the annular regions of the complex plane, and explore how the negative Kovalevskaya exponent n-n varies with the choice of region (disk, annulus, or exterior) for the expansion. We begin by rewritting our general solution (4.7) into a sum over simple poles:
$x_{1}(z)=\sum_{j=1}^{m}\frac{1}{z-z_{j}},$ (5.1)
where $z_{1},\dots,z_{m}$ are the singularities of the solution.
Next, we introduce the transformation $w:=z-z_{1}$ and $\delta_{j}:=z_{j}-z_{1}$ for $j\geq 2$, and rewrite the solution as:
$x_{1}(w)=\frac{1}{w}+\sum_{j=2}^{m}\frac{1}{w-\delta_{j}}.$ (5.2)
For each term $\frac{1}{w-\delta_{j}}$, we examine the geometric series expansions depending on the region where the expansion is performed.
Geometric series expansion
We consider the following two cases based on the relative magnitudes of $|w|$ and $|\delta_{j}|$:
  • Inside the disc ($|w|<|\delta_{j}|$): In this case, the geometric series expansion is:
    $\frac{1}{w-\delta_{j}}=-\frac{1}{\delta_{j}}\cdot\frac{1}{1-w/\delta_{j}}=-\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}},$ (5.3)
    which does not contain a $w^{-1}$ term.
  • Outside the disc ($|w|>|\delta_{j}|$): Here, the expansion becomes:
    $\frac{1}{w-\delta_{j}}=\frac{1}{w}\cdot\frac{1}{1-\delta_{j}/w}=\frac{1}{w}\sum_{k=0}^{\infty}\Big(\frac{\delta_{j}}{w}\Big)^{k},$ (5.4)
    where the leading term is $1/w$.
Annular Regions
Let $0<r_{2}\leq r_{3}\leq\cdots\leq r_{m}$ be the ordered moduli of $\delta_{j}$ (i.e. the radii $|\delta_{j}|$ in nondecreasing order). Define annular regions
$A_{n}:=\{w\in{\mathbb{C}}:\ r_{n}<|w|<r_{n+1}\},\qquad n=1,\dots,m,$ (5.5)
with the conventions $r_{1}:=0$ and $r_{m+1}:=+\infty$. Thus: $A_{1}$ is a punctured disc around $w=0$, intermediate $A_{n}$ are annuli, and $A_{m}$ is the exterior region.
For a fixed annulus $A_{n}$, split the set of poles other than $z_{1}$ into
$I_{\rm in}(n):=\{\,j\in\{2,\dots,m\}:\ |\delta_{j}|<|w|\ \text{for }w\in A_{n}\,\},\qquad I_{\rm out}(n):=\{2,\dots,m\}\setminus I_{\rm in}(n).$
By construction, $|I_{\rm in}(n)|=n-1$ and $|I_{\rm out}(n)|=m-n$.
Proposition 5.1.
Fix $n\in\{1,\dots,m\}$ and an annulus $A_{n}$ as in (5.5). Then $x_{1}(w)$ admits on $A_{n}$ a Laurent expansion of the form
$x_{1}(w)=\frac{n}{w}+H_{n}(w),$ (5.6)
where $H_{n}(w)$ is holomorphic on $A_{n}$ and admits a Laurent series that converges normally on $A_{n}$. Specifically, the explicit form of $H_{n}(w)$ is:
$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}\;-\;\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$ (5.7)
In particular, the coefficient of $w^{-1}$ in this Laurent expansion is $n$.
Proof.
We begin by recalling the expression for $x_{1}(w)$ given in (5.2). For each $j\in I_{\rm in}(n)$, we have $|\delta_{j}/w|<1$, so the series expansion (5.3) converges absolutely and uniformly on compact subsets of $A_{n}$. Similarly, for each $j\in I_{\rm out}(n)$, we have $|w/\delta_{j}|<1$, so the series (5.4) converges absolutely and uniformly on compact subsets of $A_{n}$.
Substituting these expansions into (5.2) yields:
$x_{1}(w)=\frac{1}{w}+\sum_{j\in I_{\rm in}(n)}\left(\frac{1}{w}+\sum_{k\geq 1}\frac{\delta_{j}^{k}}{w^{k+1}}\right)+\sum_{j\in I_{\rm out}(n)}\left(-\sum_{k\geq 0}\frac{w^{k}}{\delta_{j}^{k+1}}\right).$
Thus, the coefficient of $w^{-1}$ is computed as:
$1+|I_{\rm in}(n)|=1+(n-1)=n.$
The remaining terms form the Laurent series $H_{n}(w)$, which is holomorphic on $A_{n}$ and has the explicit form:
$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}-\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$
Remark 5.2.
By Proposition 5.1, the negative Kovalevskaya exponent can be interpreted as the annulus index (disk/annulus/exterior), which indicates the residue of $x_{1}$ at the movable pole. For $m\geq 3$, there are $m$ regions $A_{n}$, and thus multiple admissible Laurent expansions around a given movable pole. The inner expansion corresponds to $A_{1}$ (the punctured disk), the outer expansion corresponds to $A_{m}$ (the exterior region), and the intermediate annuli $A_{n}$ provide Laurent expansions with contributions from both inner and outer regions, with the leading coefficient being $n\in\{1,\dots,m\}$.